# Test-time Correlation Alignment

### v1

@@ Abstract
Deep neural networks often experience performance drops due to distribution shifts between training and test data. Although domain adaptation offers a solution, privacy concerns restrict access to training data in many real-world scenarios. This restriction has spurred interest in Test-Time Adaptation (TTA), which adapts models using only unlabeled test data. However, current TTA methods still face practical challenges: (1) a primary focus on instance-wise alignment, overlooking CORrelation ALignment (CORAL) due to missing source correlations; (2) complex backpropagation operations for model updating, resulting in overhead computation and (3) domain forgetting.
To address these challenges, we provide a theoretical analysis to investigate the feasibility of Test-time Correlation Alignment (TCA), demonstrating that correlation alignment between high-certainty instances and test instances can enhance test performances with a theoretical guarantee. Based on this, we propose two simple yet effective algorithms: LinearTCA and LinearTCA+. LinearTCA applies a simple linear transformation to achieve both instance and correlation alignment without additional model updates, while LinearTCA+ serves as a plug-and-play module that can easily boost existing TTA methods. Extensive experiments validate our theoretical insights and show that TCA methods significantly outperforms baselines across various tasks, benchmarks and backbones. Notably, LinearTCA improves adaptation accuracy by 5.88% on OfficeHome dataset, while using only 4% maximum GPU memory usage and 0.6% computation time compared to the best baseline TTA method. Our code is available at https://github.com/youlj109/TCA.
@@ 1 Introduction
Figure 1: An intuitive demonstration of the existing limitations. (a) Feature correlation alterations compared to the source domain, showing an increasing trend with domain shifts. (b) Computation time and maximum GPU memory usage of various TTA methods on the CIFAR-10-C dataset, where existing methods incur significant computational overhead. (c) Performance of each TTA method on the source domain after adaptation in the test domain, highlighting the difficulty in retaining source domain knowledge.
@@
However, DA methods are practically difficult when pretrained models are publicly available but the training data and training process remain inaccessible due to privacy and resource restrictions (Liang et al., 2024). To address such a source-inaccessible domain shifts task at test time, test-time adaptation (TTA) (Gong et al., 2024; Su et al., 2024a, b) has emerged as a rapidly progressing research topic. Although some recent attempts have been made to handle this task, current TTA methods still face several limitations:
@@
To address the above issues, applying “effective and frustratingly simple” CORAL in TTA seems an intuitive solution. However, the lack of access to source data makes this approach highly challenging. Consequently, we first investigate the feasibility of Test-time Correlation Alignment (TCA) by exploring two key questions: (1) Can we construct a “pseudo-source correlation” to approximate the original source correlation? (2) Can TCA based on this pseudo-source correlation enable effective TTA? We provide a theoretical analysis, showing that aligning correlations between high-certainty instances and test instances can enhance performances on test domains with a theoretical guarantee. Building on this, we propose two simple yet effective methods: LinearTCA and LinearTCA+. Specifically, we first compute the “pseudo-source correlation” by using $k$ high-certainty instances. Then, LinearTCA aligns correlation through simple linear transformations of embeddings without model updates, resulting in minimal computation and keeping source domain knowledge. While LinearTCA+ serves as a plug-and-play module that can easily boost existing TTA methods.
Main Findings and Contributions: (1) We introduce a novel and practical paradigm for TTA, termed Test-time Correlation Alignment (TCA). The construction of the pseudo-source correlation and the adaptation effectiveness are theoretically guaranteed. (2) Based on our analysis, we propose two simple yet effective methods: LinearTCA and LinearTCA+ to explore the effectiveness of TCA, as well as its potential as a plug-and-play module when combined with other TTA methods. (3) We conduct experiments to validate our theoretical insights and perform a comprehensive comparison of LinearTCA and LinearTCA+ against existing TTA methods across various benchmarks, backbones, and tasks. This evaluation encompasses multiple performance aspects, including accuracy, efficiency, and resistance to forgetting. The results demonstrate that LinearTCA achieves outstanding performance, while LinearTCA+ robustly boosts TTA methods in various conditions. (4) Further in-depth experimental analysis reveals the effective range of LinearTCA and provides valuable insights for future work.
@@ 2.2 Correlation Alignment (CORAL)
The aim of correlation alignment (CORAL) (Sun et al., 2017; Cheng et al., 2021a; Sun & Saenko, 2016; Sun et al., 2016; Das et al., 2021; Rahman et al., 2020b) is to minimize the distance of the second-order statistics (covariance) between the source and test features. Specifically, let $Z_{s}=\{z_{s}^{i}\}_{i=1}^{n_{s}}\in\mathbb{R}^{n_{s}\times d}$ denotes the feature matrix from the source domain, and $Z_{t}=\{z_{t}^{i}\}_{i=1}^{n_{s}}\in\mathbb{R}^{n_{t}\times d}$ denotes the feature matrix from the test domain. CORAL computes the covariance matrices of the source features $Z_{s}$ and test features $Z_{t}$, and aligns correlation by minimizing the Frobenius norm of their two covariance matrices. The covariance matrix is computed as below:
@@ 2.3 Problem Statement
Existing TTA methods suffer from overlooking feature correlation, overhead computation and domain forgetting. Research and practice have demonstrated that CORAL is both effective and “frustratingly easy” to implement on DA. However, due to privacy and resource constraints in TTA, it is impossible to compute the source correlation. This limitation hinders the application of CORAL in such real-world scenarios, i.e. test-time correlation alignment (TCA).
@@ 3 Theoretical Studies
In this section, we conduct an in-depth theoretical analysis of TCA based on domain adaptation and learning theory. We focus on two key questions: (1) Can we construct a “pseudo-source correlation” to approximate the original source correlation? (2) Can TCA based on this pseudo-source correlation enable effective TTA? Before discussing the main results, we first state some necessary assumptions and concepts. Missing proofs and detailed explanations are provided in Appendix B.
{mdframed}
@@ 3.1 Correlation of high-certainty test instances approximates the source correlation
Let $h_{\theta}(\cdot)=g(f(\cdot))$ be an L-Lipschitz continuous hypothesis on $\mathcal{H}$. $D_{s}$ and $D_{t}$ represent the source and test data, respectively. Let $\Omega:=\bigcup_{x\in\mathbb{D}_{t}}\mathcal{B}(x,r^{*})$ as the set of source instances near the test data, we sample $k$ instances from $\Omega$ and $\mathbb{D}_{t}$ to obtain $[X_{s},Z_{s},P_{s}]$ and $[X_{t},Z_{t},P_{t}]$ by $h_{\theta}(\cdot)$, respectively. Per 3.2, 3.3 and 3.4, with a probability of at least $1-exp(-\frac{k^{2}}{c_{t}\mu^{-}\pi_{d_{I}}(r^{*}/L)^{d_{I}}}+\log k)$, we have
@@
where $\pi_{d_{I}}=\lambda(\mathcal{B}(0,1))$ is the volume of the $d_{I}$ dimension unit ball and $d_{I}$ is the dimension of input $x$. Furthermore, considering the source correlation $\Sigma_{s}=\mathbb{E}[\tilde{Z_{s}}^{T}\tilde{Z_{s}}]$ and the test correlation $\Sigma_{t}=\tilde{Z_{t}}^{T}\tilde{Z_{t}}$, where $\tilde{Z}_{s}$ and $\tilde{Z}_{t}$ are the centered matrices. With a probability of at least $\min(1-exp(-\frac{k^{2}}{c_{t}\mu^{-}\pi_{d_{I}}(r^{*}/L)^{d_{I}}}+\log k),1-\delta)$, the correlation distance $\lVert\Sigma_{s}-\Sigma_{t}\rVert$ is bounded by:
$\lVert\Sigma_{s}-\Sigma_{t}\rVert_{F}\leq$
@@
where $\hat{Y}_{t}$ is the one-hot encoding of $P_{t}$, $A=\lVert o(kr^{*})\rVert+k\epsilon(h_{\theta}(X_{t}))+k\epsilon(h_{\theta}(X_{s}))$ represents the output error of the sampled instances, and $B=\sqrt{\frac{\log(2/\delta)}{2k}}$ is the sampling error. Theorem 3.5 implies the followings: (1) In Eq. 3.1, the terms $X_{s}$, $Z_{s}$, and $J_{g}(Z_{s})$ remain unchanged with the same source data. The primary factor influencing the correlation distance $\|\Sigma_{s}-\Sigma_{t}\|$ is prediction uncertainty $\|\hat{Y}_{t}-P_{t}\|_{F}$ and error of sampled instances $\epsilon(h_{\theta}(X_{t}))$. (2) Intuitively, previous studies (Gui et al., 2024; Niu et al., 2022; Yuan et al., 2024) empirically suggest that instances with higher output certainty have less output error. In other words, with a smaller divergence between the prediction $P_{t}$ and its one-hot encoding $\hat{Y}_{t}$, both uncertainty $\|\hat{Y}_{t}-P_{t}\|_{F}$ and error $\epsilon(h_{\theta}(X_{t}))$ will decrease, resulting in a smaller correlation distance. (3) Therefore, a reasonable pseudo-source construction method is to select the $k$ test instances with the smallest $\|\hat{Y}_{t}-P_{t}\|_{F}$ values (i.e. high-certainty test instances) and compute their correlation matrix.
@@ 4.1 Pseudo-Source
where $\omega_{t}^{k}$ represents $k$-th lowest uncertainty value in $\mathcal{M}$. Finally, the Pseudo-Source correlation can be calculated as follows:
@@
where $\hat{Z}_{\hat{n}_{s}}=\{z_{t}^{i}|z_{t}^{i}\in\mathcal{M}\}$ and $\hat{n}_{s}=\left|\mathcal{M}\right|$.
@@ 4.2 Methods
Setting $W^{T}\Sigma_{t}W=\hat{\Sigma}_{s}$ and applying eigenvalue decomposition, the closed-form solution for $W$ can be derived as  [^1] To enhance the robustness of the results, we recommend using torch’s automatic gradient descent method to mitigate potential instabilities associated with eigenvalue decomposition.:
@@
PACS OfficeHome CIFAR-10C CIFAR-100C
Method ResNet-18 ResNet-50 ViT-B/16 AVG ResNet-18 ResNet-50 ViT-B/16 AVG ResNet-18 ResNet-50 ViT-B/16 AVG ResNet-18 ResNet-50 ViT-B/16 AVG AVG
Source 81.84 84.78 87.02 84.54 62.01 67.01 76.11 68.37 50.80 50.77 71.48 57.68 31.01 34.02 51.71 38.91 62.38
BN 82.65 84.99 - - 62.05 66.30 - - 73.70 72.24 - - 48.38 48.41 - - -
TENT 85.23 88.07 84.98 86.09 63.09 67.67 76.95 69.24 75.21 72.33 71.42 73.01 50.82 50.12 52.72 51.22 69.89
EATA 83.30 84.68 86.60 84.86 62.49 67.01 76.98 68.83 73.86 72.38 73.67 73.30 49.71 49.89 62.40 54.00 70.25
SAR 85.41 85.79 87.12 86.11 62.51 67.94 76.66 69.04 73.97 73.37 71.48 72.94 51.60 50.25 54.29 52.05 70.03
TIPI 87.39 88.01 87.98 87.79 63.25 68.36 77.09 69.57 76.10 72.46 71.38 73.35 50.61 50.30 52.36 51.09 70.45
TEA 87.19 88.75 87.37 87.77 63.43 68.56 76.15 69.38 76.20 72.54 71.45 73.41 50.67 50.21 52.31 51.06 70.40
TSD 87.83 89.99 83.43 87.08 62.47 68.63 75.49 68.87 76.93 73.23 71.47 73.88 49.35 49.60 51.74 50.23 70.01
@@
Table 1: Accuracy comparison of different TTA methods based on ResNet-18/50 and ViT-B/16 backbones. The best results are highlighted in boldface, and the second ones are underlined. “-” indicates that ViT-B/16 does not include any BN layers.

### v2

@@ Abstract
Deep neural networks often degrade under distribution shifts. Although domain adaptation offers a solution, privacy constraints often prevent access to source data, making Test-Time Adaptation (TTA)—which adapts using only unlabeled test data—increasingly attractive. However, current TTA methods still face practical challenges: (1) a primary focus on instance-wise alignment, overlooking CORrelation ALignment (CORAL) due to missing source correlations; (2) complex backpropagation operations for model updating, resulting in overhead computation and (3) domain forgetting. To address these challenges, we provide a theoretical analysis to investigate the feasibility of Test-time Correlation Alignment (TCA), demonstrating that correlation alignment between high-certainty instances and test instances can enhance test performances with a theoretical guarantee. Based on this, we propose two simple yet effective algorithms: LinearTCA and LinearTCA+. LinearTCA applies a simple linear transformation to achieve both instance and correlation alignment without additional model updates, while LinearTCA+ serves as a plug-and-play module that can easily boost existing TTA methods. Extensive experiments validate our theoretical insights and show that TCA methods significantly outperforms baselines across various tasks, benchmarks and backbones. Notably, LinearTCA achieves higher accuracy with only 4% GPU memory and 0.6% computation time compared to the best TTA baseline. It also outperforms existing methods on CLIP over 1.86%. Code: https://github.com/youlj109/TCA.
@@
Figure 1: Illustration of key limitations in existing TTA methods. (a) Correlation distance increases with domain shifts. (b) Computation time and peak GPU memory usage on CIFAR-10-C, showing high overhead of existing methods. (c) Source domain performance after test-time adaptation, revealing challenges in retaining source knowledge.
@@ 1 Introduction
However, DA methods are practically difficult when pretrained models are publicly available but the training data and training process remain inaccessible due to privacy and resource restrictions (Liang et al., 2024). To address such a source-inaccessible domain shifts task at test time, test-time adaptation (TTA) (Gong et al., 2024; Su et al., 2024a, b; You et al., 2025) has emerged as a rapidly progressing research topic. Although some recent attempts have been made to handle this task, current TTA methods still face several limitations:
@@
To address the above challenges, applying the “effective and frustratingly simple” CORAL method to TTA appears intuitive—but the lack of source data makes it highly challenging. We thus explore the feasibility of Test-time Correlation Alignment (TCA) by posing key questions: (1)Can we construct a pseudo-source correlation that approximates the true source correlation? (2) Can this enable effective TTA? We provide a theoretical analysis showing that aligning correlations between high-certainty and test instances improves test-time performance with guarantees. Based on this, we propose two simple yet effective methods: LinearTCA and LinearTCA+. Specifically, we first compute the “pseudo-source correlation” by using $k$ high-certainty instances. Then, LinearTCA aligns correlation through simple linear transformations of embeddings without model updates, resulting in minimal computation and keeping source domain knowledge. While LinearTCA+ serves as a plug-and-play module that can easily boost existing TTA methods.
Main Findings and Contributions: (1) We introduce a novel and practical paradigm for TTA, termed Test-time Correlation Alignment (TCA). The construction of the pseudo-source correlation and the adaptation effectiveness are theoretically guaranteed. (2) Based on our analysis, we develop two simple yet effective methods—LinearTCA and LinearTCA+—to validate TCA’s effectiveness and its plug-and-play potential with other TTA approaches. (3)We conduct comprehensive experiments to validate our theoretical insights and compare performance across diverse benchmarks, backbones, and tasks, evaluating accuracy, efficiency, and resistance to forgetting. Results show that LinearTCA achieves outstanding performance, while LinearTCA+ robustly boosts other TTA methods under various conditions. (4) Further in-depth experimental analysis reveals the effective range of LinearTCA and provides valuable insights for future work.
@@ 2.2 Correlation Alignment (CORAL)
The aim of correlation alignment (CORAL) (Sun et al., 2017; Cheng et al., 2021a; Sun & Saenko, 2016; Sun et al., 2016; Das et al., 2021; Rahman et al., 2020b) is to minimize the distance of the second-order statistics (covariance) between the source and test features. Specifically, let $Z_{s}=\{z_{s}^{i}\}_{i=1}^{n_{s}}\in\mathbb{R}^{n_{s}\times d}$ denotes the feature matrix from the source domain, and $Z_{t}=\{z_{t}^{i}\}_{i=1}^{n_{t}}\in\mathbb{R}^{n_{t}\times d}$ denotes the feature matrix from the test domain. CORAL computes the covariance matrices of the source features $Z_{s}$ and test features $Z_{t}$, and aligns correlation by minimizing the Frobenius norm of their two covariance matrices. The covariance matrix is computed as below:
@@ 2.3 Problem Statement
Existing TTA methods suffer from overlooking feature correlation, overhead computation and domain forgetting. Research and practice have demonstrated that CORAL is both effective and “frustratingly easy” to implement on DA. Since TTA is a subfield of DA, it is a natural extension to apply CORAL within TTA frameworks. However, due to privacy and resource constraints in TTA, it is impossible to compute the source correlation. This limitation hinders the application of CORAL in such real-world scenarios, i.e. test-time correlation alignment (TCA).
@@ 3 Theoretical Studies
In this section, we conduct an in-depth theoretical analysis of TCA based on domain adaptation and learning theory. We focus on two key questions: (1) Can we construct a “pseudo-source correlation” to approximate the original source correlation? (2) Can TCA based on this pseudo-source correlation enable effective TTA? Before discussing the main results, we first state some necessary assumptions and concepts. Missing proofs and detailed explanations are provided in Appendix B. {mdframed}[backgroundcolor=gray!10, topline=false, bottomline=false, leftline=false, rightline=false, innertopmargin=1pt, innerbottommargin=1pt]
@@ 3.1 Correlation of high-certainty test instances approximates the source correlation
Let $h_{\theta}(\cdot)=g(f(\cdot))$ be an L-Lipschitz continuous hypothesis on $\mathcal{H}$. $\Omega:=\bigcup_{x\in\mathbb{D}_{t}}\mathcal{B}(x,r^{*})$ is the set of balls near the test data. We sample $k$ source instances from $\mathbb{D}_{s}\cap\Omega$ and $k$ test instances from $\mathbb{D}_{t}$ to obtain $[X_{s},Z_{s},P_{s}]$ and $[X_{t},Z_{t},P_{t}]$ by $h_{\theta}(\cdot)$, respectively. Per 3.2, 3.3 and 3.4, with a probability of at least $1-\exp\left(-\frac{(c_{t}\mu^{-}\pi_{d_{I}}r^{d_{I}}n_{s}-1)^{2}}{2c_{t}\mu^{-}\pi_{d_{I}}r^{d_{I}}n_{s}}+\log k\right)$, we have
@@
where $\pi_{d_{I}}=\lambda(\mathcal{B}(0,1))$ is the volume of the $d_{I}$ dimension unit ball and $d_{I}$ is the dimension of input $x$. Furthermore, considering the true source correlation $\Sigma_{s}=\mathbb{E}[\tilde{Z_{s}}^{T}\tilde{Z_{s}}]$ and the pseudo-source correlation $\hat{\Sigma}_{s}=\tilde{Z_{t}}^{T}\tilde{Z_{t}}$, where $\tilde{Z}_{s}$ and $\tilde{Z}_{t}$ are centered. With a probability of at least $\min(1-\exp\left(-\frac{(c_{t}\mu^{-}\pi_{d_{I}}r^{d_{I}}n_{s}-1)^{2}}{2c_{t}\mu^{-}\pi_{d_{I}}r^{d_{I}}n_{s}}+\log k\right),1-\delta)$, the correlation distance $\lVert\Sigma_{s}-\hat{\Sigma}_{s}\rVert$ is bounded by:
$\lVert\Sigma_{s}-\hat{\Sigma}_{s}\rVert_{F}\leq$
@@
where $\hat{Y}_{t}$ is the one-hot encoding of $P_{t}$, $A=\lVert o(kr^{*})\rVert+k\epsilon(h_{\theta}(X_{t}))+k\epsilon(h_{\theta}(X_{s}))$ represents the output error of the sampled instances, and $B=\sqrt{\frac{\log(2/\delta)}{2k}}$ is the sampling error. Theorem 3.5 implies the followings: (1) In Eq. (3.1), the terms $X_{s}$, $Z_{s}$, and $J_{g}(Z_{s})$ remain unchanged with the same source data. The primary factor influencing the correlation distance $\|\Sigma_{s}-\hat{\Sigma}_{s}\|$ is prediction uncertainty $\|\hat{Y}_{t}-P_{t}\|_{F}$ and output error of the sampled instances $\epsilon(h_{\theta}(X_{t}))$. (2) Intuitively, previous studies (Gui et al., 2024; Niu et al., 2022; Yuan et al., 2024) empirically suggest that instances with higher output certainty have less output error. In other words, with a smaller divergence between the prediction $P_{t}$ and its one-hot encoding $\hat{Y}_{t}$, both uncertainty $\|\hat{Y}_{t}-P_{t}\|_{F}$ and error $\epsilon(h_{\theta}(X_{t}))$ will decrease, resulting in a smaller correlation distance. (3) Therefore, a reasonable pseudo-source construction method is to select the $k$ test instances with the smallest $\|\hat{Y}_{t}-P_{t}\|_{F}$ values (i.e. high-certainty test instances) and compute their correlation matrix as pseudo-source correlation.
@@ 4.1 Pseudo-Source
where $\omega_{min}^{k}$ represents $k$-th lowest uncertainty value in $\mathcal{M}$. Finally, the Pseudo-Source correlation can be calculated as follows:
@@
where $\hat{Z}_{s}=\{z_{t}^{i}|z_{t}^{i}\in\mathcal{M}\}$ and $\hat{n}_{s}=\left|\mathcal{M}\right|$.
@@ 4.2 Methods
Setting $W^{T}\Sigma_{t}W=\hat{\Sigma}_{s}$ and applying eigenvalue decomposition, the closed-form solution for $W$ can be derived as  [^1] To enhance the robustness of the results, we recommend using torch’s automatic gradient descent method to mitigate potential instabilities associated with eigenvalue decomposition. For the following experiments, we implement this method with a fixed learning rate of 1e-3.:
@@
Domain PACS OfficeHome DomainNet
Method ResNet-18 ResNet-50 ViT-B/16 AVG ResNet-18 ResNet-50 ViT-B/16 AVG ResNet-18 ResNet-50 ViT-B/16 AVG AVG
SOURCE 81.84 84.78 87.02 84.54 62.01 67.01 76.11 68.37 39.13 43.58 50.29 44.33 65.75
BP-Free BN 82.65 84.99 - - 62.05 66.30 - - 37.93 41.94 - - -
T3A 83.20 85.71 88.06 85.66 63.26 67.85 78.87 69.99 40.62 44.92 53.94 46.49 67.38
AdaNPC 83.48 86.12 89.11 86.24 62.88 67.05 77.26 69.07 40.50 45.17 53.28 46.32 67.21
BP-Based TENT 85.23 88.07 84.98 86.09 63.09 67.67 76.95 69.24 39.42 43.97 39.96 41.12 65.48
PLC 83.16 86.59 87.97 85.91 62.22 66.44 76.51 68.39 37.96 41.63 47.29 42.30 65.53
EATA 83.30 84.68 86.60 84.86 62.49 67.01 76.98 68.83 41.65 46.89 54.40 47.65 67.11
SAR 85.41 85.79 87.12 86.11 62.51 67.94 76.66 69.04 38.49 42.19 42.81 41.16 65.44
TIPI 87.39 88.01 87.98 87.79 63.25 68.36 77.09 69.57 36.05 44.08 39.70 39.94 65.77
TEA 87.19 88.75 87.37 87.77 63.43 68.56 76.15 69.38 39.43 43.48 48.41 43.78 66.98
TSD 87.83 89.99 83.43 87.08 62.47 68.63 75.49 68.87 38.59 42.12 48.72 43.14 66.36
@@
ImgCop CIFAR-10-C CIFAR-100-C ImageNet-C
Method ResNet-18 ResNet-50 ViT-B/16 AVG ResNet-18 ResNet-50 ViT-B/16 AVG ResNet-18 ResNet-50 ViT-B/16 AVG AVG
SOURCE 50.80 50.77 71.48 57.68 31.01 34.02 51.71 38.91 14.70 18.15 39.83 24.23 40.27
BP-Free BN 73.70 72.24 - - 48.38 48.41 - - 27.59 32.06 - - -
T3A 58.89 54.87 74.21 62.65 32.52 34.94 54.24 40.57 14.56 18.05 39.78 24.13 42.45
AdaNPC 57.72 54.75 74.60 62.36 29.70 32.27 53.21 38.39 11.93 15.62 36.78 21.44 40.73
BP-Based TENT 75.21 72.33 71.48 73.01 50.82 50.12 52.72 51.22 35.39 41.32 48.01 41.57 55.27
PLC 73.72 72.34 71.46 72.51 48.35 48.38 51.71 49.48 27.59 32.06 38.74 32.80 51.59
EATA 73.86 72.38 73.67 73.30 49.71 49.89 62.40 54.00 39.19 48.17 64.36 50.58 59.29
SAR 73.97 73.37 71.48 72.94 51.60 50.25 54.29 52.05 38.55 46.30 57.94 47.60 57.53
TIPI 76.10 72.46 71.48 73.35 50.61 50.30 52.36 51.09 35.73 41.87 48.50 42.03 55.49
TEA 76.20 72.54 71.48 73.41 50.67 50.21 52.31 51.06 32.38 38.90 41.37 37.55 54.01
TSD 76.93 73.23 71.47 73.88 49.35 49.60 51.74 50.23 30.11 35.08 41.33 35.51 53.20
[…]
Table 1: Accuracy comparison of different TTA methods based on ResNet-18/50 and ViT-B/16 backbones. The upper part of the table corresponds to the domain generalization task, while the lower part corresponds to the image corruption task. The best results are highlighted in boldface, and the second ones are underlined. “-” indicates that ViT-B/16 does not include any BN layers.